# Searching for leptophilic composite asymmetric dark sector at $e^{+}e^{-}$ colliders

### v1

@@ IV.2 Detecting dark showers at the Belle II
Unlike most light DM searches at the Belle II, the mono-photon search strategy lacks sufficient discriminating power between signal and background events in this model. Although undetectable dark pions produce missing energy signatures, their kinematic distributions remain similar to SM backgrounds. Because the mediator mass exceeds the center-of-mass energy at the Belle II, this absence of light mediator production results in smooth, continuous photon energy distributions that resemble SM backgrounds Bernreuther:2022jlj . We therefore exclude mono-photon constraints from Belle II in this analysis.
@@ IV.3 Long-lived dark pion searches in proposed GAZELLE detector
We apply Displaced Decay Counter (DDC) Domingo:2023dew for the L-GAZELLE detector simulation [^4] The L-GAZELLE detector is approximated with LG-B1 (single-layer approximation within $8.5\%$ error) and LG-B2 (four-layer approximation within $5.0\%$ error).. The decay probability of LLPs within a certain range is calculated as: $P(\ell_{1},\ell_{2})=\exp\left(-\frac{\ell_{1}}{\gamma|\vec{v}|\tau_{0}}\right)-\exp\left(-\frac{\ell_{2}}{\gamma|\vec{v}|\tau_{0}}\right)$, where $\gamma\equiv(1-|\vec{v}|^{2})^{-1/2}$, $|\vec{v}|$ and $\tau_{0}$ are the velocity and proper decay length of a LLP with the natural units. The weight for each layer is $\Delta\phi/(2\pi)$, where $\Delta\phi$ is the angular aperture of each layer. The weighted sum of probabilities from all detector layers gives the acceptance $A_{i}$ for each LLPi, which represents the detection probability of each LLP. For event selections, we calculate the probabilities of detecting at least one LLP and at least two LLPs:
  • At least one LLP detected: $\langle P(\geq 1)\rangle=1-\langle\prod(1-A_{i})\rangle$;
@@
For the signal efficiency $\epsilon_{S}$ obtained for both scenarios, the expected signal yield is given by $N_{S}=\mathcal{L}\times\sigma(e^{+}e^{-}\to\bar{q}_{d}q_{d})\times\epsilon_{S}$, assuming an optimistic integrated luminosity $\mathcal{L}=50~\text{ab}^{-1}$. The $90\%$ C.L. exclusion contours correspond to the lower boundaries, shown as orange-red (at least one LLP detected) and brown (at least two LLPs detected) solid lines in Fig. 6. Compared to recast BaBar and projected Belle II constraints, the far-detector L-GAZELLE coverage extends to smaller $\kappa$ and larger $m_{X}$ values.
@@
Table 5: Detection efficiency $\langle P(\geq 1)\rangle$ and $\langle P(\geq 2)\rangle$ for various dark pion proper decay length $\tau_{0}$ and dark pion mass $m_{\pi_{d}}$ values. Corresponding production cross-sections are $7.15\times 10^{-2}$ fb for $m_{\pi_{d}}=0.32$ GeV at $\tau_{0}=54.91$ m, $6.85\times 10^{-2}$ fb for $m_{\pi_{d}}=0.7$ GeV at $\tau_{0}=5.25$ m, and $6.45\times 10^{-2}$ fb for $m_{\pi_{d}}=1$ GeV at $\tau_{0}=1.80$ m.
@@
We further evaluated detection efficiencies for different $m_{\pi_{d}}$ and $\tau_{0}$ values, as shown in Table 5. The detection efficiency for at least two LLPs ($\langle P(\geq 2)\rangle$) is significantly lower than for at least one LLP ($\langle P(\geq 1)\rangle$). For BP1, $\langle P(\geq 2)\rangle$ is about two orders of magnitude smaller than $\langle P(\geq 1)\rangle$. As the $m_{\pi_{d}}$ increases, the efficiency gap widens, particularly at $m_{\pi_{d}}=1$ GeV, where events typically produce two LLPs that travel nearly back-to-back, making it highly unlikely for both to be detected within L-GAZELLE. Since $\tau_{0}$ and the production cross-section are anti-correlated, a decrease in $\tau_{0}$ enhances the cross-section but causes a more rapid decline in detection efficiency. This efficiency loss prevents meaningful constraints for BP2 and BP3 in the scenario requiring at least two detected LLPs. Consequently, the brown solid line corresponding to BP2 and BP3 is absent in Fig. 6.
V Dense dark showers at the CEPC
@@
This section further investigates the detection prospects at a high-energy $e^{+}e^{-}$ collider, focusing on the proposed circular electron positron collider (CEPC) with a center-of-mass energy of $\sqrt{s}=240~\text{GeV}$ and an integrated luminosity of $5.6~\text{ab}^{-1}$. We also study the exclusion limits for two mass ranges of the dark pion: $10~\text{MeV}\leq m_{\pi_{d}}<2m_{\mu}$, and $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, which lead to the mono-photon and displaced muon jet (DMJ) signatures.
V.1 The mono-photon signature
@@
Figure 9: Exclusion regions for different dark pion mass ranges. The first panel shows the mono-photon signature for $10~\text{MeV}<m_{\pi_{d}}<2m_{\mu}$, while the remaining three panels show benchmark points $m_{\pi_{d}}=0.32$, $0.7$ and $1~\text{GeV}$ for both CEPC mono-photon and DMJ signatures, covering $2m_{\mu}<m_{\pi_{d}}<2m_{\tau}$.
@@
For the dark pion mass range $10~\text{MeV}\leq m_{\pi_{d}}<2m_{\mu}$, we maintain the conservative setting $m_{\pi_{d}}=f_{\pi_{d}}=0.2~\text{GeV}$ and impose the requirement that the $\pi_{d}$ proper decay length exceeds $6.08~\text{m}$ to guarantee that all dark mesons decay outside the CEPC detector volume. The resulting upper bounds on $\kappa$, shown as the yellow dashed curve in Fig. 9, reach values of $\mathcal{O}(1-10)$. We recast the previous CEPC mono-photon analysis, derived for the $t$-channel scalar mediator effective operator with cutoff scale $\Lambda_{t}$ (corresponding to the $t$-channel case in Eq. (2.3)) and the upper right panel (H-mode) of Fig. 16 in Ref. Liu:2019ogn , to our leptophilic composite ADM framework through the relation:
@@
Assuming light dark quarks and determining the lower limits on $\Lambda$, the yellow region in Fig. 9 indicates parameter space where the predicted cross-sections exceed the experimental bounds.
@@ V.2 The displaced muon jet signature
For the dark pion mass range $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, unlike low-energy $e^{+}e^{-}$ colliders, due to the high center-of-mass energy, the initially produced dark quarks undergo dark showering and hadronization processes, yielding a large number of dark pions with the same model parameter settings. These dark pions subsequently decay to muons, resulting in a high multiplicity of displaced muons that collectively form displaced muon jets (DMJs).
@@
Hadron Calorimeter $2.30\text{\,}\mathrm{m}$ $3.34\text{\,}\mathrm{m}$ $30.00\text{\,}\mathrm{m}\mathrm{m}$
Muon System $4.40\text{\,}\mathrm{m}$ $6.08\text{\,}\mathrm{m}$ $20.00\text{\,}\mathrm{m}\mathrm{m}$
@@
The detailed definition of DMJ in this analysis is as follows. All candidate muons inside the DMJ are selected by requiring $p_{T}(\mu)>1~\text{GeV}$ and $|\eta_{\mu}|<3.0$, parent particles ($\pi_{d}$’s) with $R_{xy}$ within one of the four regions defined in Table 6, and muon pairs with $\Delta\theta>0.02$ radians. Muons satisfying these criteria are identified as displaced muons. We employ the anti-$k_{t}$ algorithm Cacciari:2008gp as implemented in FASTJET Cacciari:2011ma to cluster these muons into jets, referred to as DMJs, using a jet radius parameter of $R=0.5$. In our analysis, we require at least six muons per jet, which renders the DMJ searches effectively background-free Zhang:2021orr .
Figure 10: The main kinematic distributions for the DMJ analysis at the CEPC with three benchmark points, $m_{\pi_{d}}=0.32$, $0.7$, and $1$ GeV, corresponding to $\kappa=5$ and $m_{X}=1$ TeV. Top-left: The dark pion transverse decay length, $R_{xy}$; Top-right: Muon transverse momentum, $p_{T}(\mu)$; Bottom-left: The number of muons in each muon jet; Bottom-right: The opening angle $\Delta\theta$ between two muons originating from the same dark pion.
We then select the leading $p_{T}$ DMJ in each event and present in Fig. 10 the main kinematic distributions of its constituent muons and their mother dark pions, using three BPs with $m_{\pi_{d}}=0.32~\text{GeV}$, $0.7~\text{GeV}$, and $1~\text{GeV}$, corresponding to $\kappa=5$ and $m_{X}=1~\text{TeV}$. The $R_{xy}$ (top-left) and $p_{T}(\mu)$ (top-right) distributions follow trends similar to those observed at low-energy $e^{+}e^{-}$ colliders, but at the CEPC they extend to significantly higher values due to its much larger collision energy. Since the CEPC detector is segmented into four non-contiguous radial regions as shown in Table 6, the $R_{xy}$ distribution reflects this structure; for three BPs considered here, most dark pions are concentrated in the innermost region, namely the vertex detector. The calculation of $R_{xy}$ also includes Gaussian smearing effects, with the transverse position resolution for displaced vertices given by CEPCStudyGroup:2018ghi :
@@
where $a=5$, $b=$10\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$\$, $p_{T}$ is the average transverse momentum of the track, and $\theta$ is its polar angle. The distribution of the number of muons inside the leading $p_{T}$ DMJ (bottom-left) illustrates the multiplicity of displaced muons. Across all three BPs, DMJs exhibit high-multiplicity features, with lighter $\pi_{d}$’s showing a multiplicity advantage. Such high-multiplicity signatures are distinctive and very difficult for SM backgrounds to mimic Zhang:2021orr . The opening angle $\Delta\theta$ between two muons originating from the same dark pion (bottom-right) increases from BP1 to BP3. This wider separation for heavier $\pi_{d}$ results from the larger $p_{T}$ carried by the muons. However, due to the finite spatial resolution of the detector, muon pairs with excessively small $\Delta\theta$ cannot be resolved and are therefore excluded from the $\pi_{d}$ reconstruction.
To further identify the signal signature, all events containing at least two DMJs, each comprising more than six displaced muons, are selected. The signal efficiency $\epsilon_{S}$ is then evaluated from the above event selection criteria. The expected signal yield is given by $N_{S}=\mathcal{L}\times\sigma(e^{+}e^{-}\to q_{d}\bar{q}_{d})\times\epsilon_{S}$. Under the background-free hypothesis, the $2\sigma$ exclusion limit is set at $N_{S}=3$, from which we derive the exclusion contour in the $(m_{X},\kappa)$ plane, as shown in Fig. 9.
@@ VI Summary and discussion
We explored novel signatures of a leptophilic QCD-like composite asymmetric dark sector at $e^{+}e^{-}$ colliders by simulating dark parton showering, hadronization, and detector response to map the experimental reach. The study reveals a sharp transition in the dark pion proper decay length at the muon threshold (Fig. 2). Recast constraints from mono-photon searches and from slepton/stau reinterpretations limit the low-mass bi-charged scalar mediator region (Figs. 3 and 4), while combined projections for BaBar, Belle II, L-GAZELLE, and CEPC produce the characteristic band-shaped sensitivity in the $(m_{X},\kappa)$ plane as summarized in Fig. 11. After applying cosmological constraints on the $\pi_{d}$ lifetime, nearly all parameter space for $10~\text{MeV}\leq m_{\pi_{d}}<2m_{\mu}$ and $m_{X}\lesssim 10~\text{TeV}$ is excluded. Additionally, for $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, a significant fraction of parameter space with $\kappa\gtrsim 1$ and $m_{X}\lesssim 10~\text{TeV}$ is also excluded.
@@
Figure 11: Combined exclusion limits for different dark pion mass ranges. The first panel shows the exclusion region for $10~\text{MeV}<m_{\pi_{d}}<2m_{\mu}$, while the remaining three panels show benchmark points $m_{\pi_{d}}=0.32$, $0.7$ and $1~\text{GeV}$ from LHC, mono-photon searches at LEP and CEPC, displaced vertex searches (BaBar, Belle II, L-GAZELLE) and DMJ search at CEPC. All four panels include the parameter space where the proper lifetime of $\pi_{d}$ exceeds 1 sec, conservatively evaluated at $m_{\pi_{d}}=0.2~\mathrm{GeV}$ for the light mass range.
@@
The primary experimental targets we identify are: (i) sparse displaced dimuon vertices at B-factories (kinematics in Figs. 57 and 8; trigger and event selection conditions are summarized in Tables 2 and 3), which Belle II can probe with realistic triggers and event selections; (ii) far detectors (L-GAZELLE) that extend sensitivity to substantially longer lifetimes; and (iii) displaced muon jet (DMJ) searches at high-energy $e^{+}e^{-}$ colliders (main kinematics shown in Fig. 10 and detector assumptions in Table 6). Although our projections adopt simplifying assumptions, lepton-flavor universality (LFU) of the mediator coupling, the benchmark relation $f_{\pi_{d}}=m_{\pi_{d}}$, and optimistic/background-free hypotheses in some regions, they nevertheless provide concrete, experimentally accessible targets and strongly motivate a coordinated program across B-factories, far detectors, and future $e^{+}e^{-}$ colliders.
@@
Finally, we outline promising directions for future work. High-energy muon colliders could directly produce the bi-charged scalar mediator via Drell–Yan processes and thereby provide an independent probe of the parameter space. Advanced data-scouting or trigger-level analyses for low-$p_{T}$ displaced dimuon pairs at CMS and LHCb could significantly strengthen LHC constraints LHCb:2020ysn ; CMS:2023hwl ; Born:2023vll . When the dark pion is heavier than $2m_{\tau}$, dark showers become $\tau$-rich; this regime will require dedicated strategies (e.g., displaced and boosted $\tau$-tagging, and optimized reconstruction of hadronic $\tau$ decays) to retain sensitivity.

### v2

@@ IV.2 Detecting dark showers at the Belle II
Unlike most light DM searches at the Belle II, the monophoton search strategy lacks sufficient discriminating power between signal and background events in this model. Although undetectable dark pions produce missing energy signatures, their kinematic distributions remain similar to SM backgrounds. Because the mediator mass exceeds the center-of-mass energy at the Belle II, this absence of light mediator production results in smooth, continuous photon energy distributions that resemble SM backgrounds Bernreuther:2022jlj . We, therefore, exclude monophoton constraints from Belle II in this analysis.
@@ IV.3 Long-lived dark pion searches in proposed GAZELLE detector
We apply a DDC Domingo:2023dew for the L-GAZELLE detector simulation. [^4] The L-GAZELLE detector is approximated with LG-B1 (single-layer approximation within $8.5\%$ error) and LG-B2 (four-layer approximation within $5.0\%$ error). The decay probability of LLPs within a certain range is calculated as: $P(\ell_{1},\ell_{2})=\exp\left(-\frac{\ell_{1}}{\gamma|\vec{v}|\tau_{0}}\right)-\exp\left(-\frac{\ell_{2}}{\gamma|\vec{v}|\tau_{0}}\right)$, where $\gamma\equiv(1-|\vec{v}|^{2})^{-1/2}$, $|\vec{v}|$ and $\tau_{0}$ are the velocity and proper decay length of a LLP with the natural units. The weight for each layer is $\Delta\phi/(2\pi)$, where $\Delta\phi$ is the angular aperture of each layer. The weighted sum of probabilities from all detector layers gives the acceptance $A_{i}$ for each LLPi, which represents the detection probability of each LLP. For event selections, we calculate the probabilities of detecting at least one LLP and at least two LLPs:
  • At least one LLP detected: $\langle P(\geq 1)\rangle=1-\langle\prod(1-A_{i})\rangle$.
@@
For the signal efficiency $\epsilon_{S}$ obtained for both scenarios, the expected signal yield is given by $N_{S}=\mathcal{L}\times\sigma(e^{+}e^{-}\to\bar{q}_{d}q_{d})\times\epsilon_{S}$, assuming an optimistic integrated luminosity $\mathcal{L}=50~\text{ab}^{-1}$. The $90\%$ C.L. exclusion contours correspond to the lower boundaries, shown as orange-red (at least one LLP detected) and brown (at least two LLPs detected) solid lines in Fig. 6. Compared to recast BaBar and projected Belle II constraints, the far detector L-GAZELLE coverage extends to smaller $\kappa$ and larger $m_{X}$ values.
@@
Table 5: Detection efficiency $\langle P(\geq 1)\rangle$ and $\langle P(\geq 2)\rangle$ for various dark pion proper decay length $\tau_{0}$ and dark pion mass $m_{\pi_{d}}$ values. Corresponding production cross sections are $5.40\times 10^{-2}$ fb for $m_{\pi_{d}}=0.32$ GeV at $\tau_{0}=54.91$ m, $6.59\times 10^{-2}$ fb for $m_{\pi_{d}}=0.7$ GeV at $\tau_{0}=5.25$ m, and $6.32\times 10^{-2}$ fb for $m_{\pi_{d}}=1$ GeV at $\tau_{0}=1.80$ m.
@@
We further evaluated detection efficiencies for different $m_{\pi_{d}}$ and $\tau_{0}$ values, as shown in Table 5. The detection efficiency for at least two LLPs $\langle P(\geq 2)\rangle$] is significantly lower than for at least one LLP $\langle P(\geq 1)\rangle$]. For BP1, $\langle P(\geq 2)\rangle$ is about 2 orders of magnitude smaller than $\langle P(\geq 1)\rangle$. As the $m_{\pi_{d}}$ increases, the efficiency gap widens, particularly at $m_{\pi_{d}}=1$ GeV, where events typically produce two LLPs that travel nearly back-to-back, making it highly unlikely for both to be detected within L-GAZELLE. Since $\tau_{0}$ and the production cross section are anticorrelated, a decrease in $\tau_{0}$ enhances the cross section but causes a more rapid decline in detection efficiency. This efficiency loss prevents meaningful constraints for BP2 and BP3 in the scenario requiring at least two detected LLPs. Consequently, the brown solid line corresponding to BP2 and BP3 is absent in Fig. 6.
V Dense dark showers at the CEPC/FCC-ee
@@
This section further investigates the detection prospects at high-energy $e^{+}e^{-}$ colliders, specifically the proposed CEPC and the FCC-ee. Both are considered to operate at a center-of-mass energy of $\sqrt{s}=240~\text{GeV}$, with integrated luminosities of $5.6~\text{ab}^{-1}$ and $5.0~\text{ab}^{-1}$, respectively. Given the expected comparable sensitivity, the analysis in the main text focuses on CEPC, while a direct comparison of the exclusion regions for FCC-ee and CEPC is provided in Appendix A. We also present exclusion limits for two distinct dark pion mass ranges: $10~\text{MeV}\leq m_{\pi_{d}}<2m_{\mu}$, and $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, which correspond to the monophoton and displaced muon jet (DMJ) signatures, respectively.
V.1 The monophoton signature
@@
Figure 9: Exclusion regions for different dark pion mass ranges. The first panel shows the monophoton signature for $10~\text{MeV}<m_{\pi_{d}}<2m_{\mu}$, while the remaining three panels show benchmark points $m_{\pi_{d}}=0.32$, $0.7$ and $1~\text{GeV}$ for both CEPC monophoton and DMJ signatures, covering $2m_{\mu}<m_{\pi_{d}}<2m_{\tau}$.
@@
For the dark pion mass range $10~\text{MeV}\leq m_{\pi_{d}}<2m_{\mu}$, we maintain the conservative setting $m_{\pi_{d}}=f_{\pi_{d}}=0.2~\text{GeV}$ and impose the requirement that the $\pi_{d}$ proper decay length exceeds $6.08~\text{m}$ to guarantee that all dark mesons decay outside the CEPC detector volume. The resulting upper bounds on $\kappa$, shown as the yellow dashed curve in Fig. 9, reach values of $\mathcal{O}(1-10)$. We recast the previous CEPC monophoton analysis, derived for the $t$-channel scalar mediator effective operator with cutoff scale $\Lambda_{t}$ [corresponding to the $t$-channel case in Eq. (2.3)] and the upper-right panel (H-mode) of Fig. 16 in Ref. Liu:2019ogn , to our leptophilic composite ADM framework through the relation
@@
Assuming light dark quarks and determining the lower limits on $\Lambda$, the yellow region in Fig. 9 indicates parameter space where the predicted cross sections exceed the experimental bounds.
@@ V.2 The displaced muon jet signature
For the dark pion mass range $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, unlike low-energy $e^{+}e^{-}$ colliders, due to the high center-of-mass energy, the initially produced dark quarks undergo dark showering and hadronization processes, yielding a large number of dark pions with the same model parameter settings. These dark pions subsequently decay to muons, resulting in a high multiplicity of displaced muons that collectively form DMJs.
@@
Hadron Calorimeter $2.30\text{\,}\mathrm{m}$ $3.34\text{\,}\mathrm{m}$ $30\text{\,}\mathrm{m}\mathrm{m}$
Muon System $4.40\text{\,}\mathrm{m}$ $6.08\text{\,}\mathrm{m}$ $20\text{\,}\mathrm{m}\mathrm{m}$
@@
The detailed definition of DMJs in this analysis is as follows. All candidate muons inside the DMJ are selected by requiring $p_{T}(\mu)>1~\text{GeV}$ and $|\eta_{\mu}|<3.0$, parent particles ($\pi_{d}$’s) with $R_{xy}$ within one of the four regions defined in Table 6, and muon pairs with $\Delta\theta>0.02$ radians. Muons satisfying these criteria are identified as displaced muons. We employ the anti-$k_{t}$ algorithm Cacciari:2008gp as implemented in FASTJET Cacciari:2011ma with a jet radius parameter of $R=0.5$ to cluster muons into jets, and subsequently identify those containing at least six muons as DMJs.
Figure 10: The main kinematic distributions for the DMJ analysis at the CEPC with three benchmark points, $m_{\pi_{d}}=0.32$, $0.7$, and $1$ GeV, corresponding to $\kappa=5$ and $m_{X}=1$ TeV. Top left: the dark pion transverse decay length, $R_{xy}$. Top right: muon transverse momentum, $p_{T}(\mu)$. Bottom left: the number of muons in each muon jet. Bottom right: the opening angle $\Delta\theta$ between two muons originating from the same dark pion.
We then select the leading $p_{T}$ DMJ in each event and present in Fig. 10 the main kinematic distributions of its constituent muons and their mother dark pions, using three BPs with $m_{\pi_{d}}=0.32~\text{GeV}$, $0.7~\text{GeV}$, and $1~\text{GeV}$, corresponding to $\kappa=5$ and $m_{X}=1~\text{TeV}$. The $R_{xy}$ (top left) and $p_{T}(\mu)$ (top right) distributions follow trends similar to those observed at low-energy $e^{+}e^{-}$ colliders, but at the CEPC they extend to significantly higher values due to its much larger collision energy. Since the CEPC detector is segmented into four noncontiguous radial regions, as shown in Table 6, the $R_{xy}$ distribution reflects this structure. For three BPs considered here, most dark pions are concentrated in the innermost region, namely the vertex detector. The calculation of $R_{xy}$ also includes Gaussian smearing effects, with the transverse position resolution for displaced vertices given by CEPCStudyGroup:2018ghi
@@
where $a=5$, $b=$10\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$\$, $p_{T}$ is the average transverse momentum of the track, and $\theta$ is its polar angle. The distribution of the number of muons inside the leading $p_{T}$ DMJ (bottom left) illustrates the multiplicity of displaced muons. Across all three BPs, DMJs exhibit high-multiplicity features, with lighter $\pi_{d}$’s showing a multiplicity advantage. Such high-multiplicity signatures are distinctive and very difficult for SM backgrounds to mimic Zhang:2021orr . The opening angle $\Delta\theta$ between two muons originating from the same dark pion (bottom right) increases from BP1 to BP3. This wider separation for heavier $\pi_{d}$ results from the larger $p_{T}$ carried by the muons. However, due to the finite spatial resolution of the detector, muon pairs with excessively small $\Delta\theta$ cannot be resolved and are, therefore, excluded from the $\pi_{d}$ reconstruction.
To further identify the signal signature, all events containing at least two DMJs are selected. The CEPC muon system comprises eight sensitive layers alternating with iron absorber plates. With a total iron thickness equivalent to 6.7 interaction lengths, this configuration offers sufficient protection against punch-through backgrounds. The detector achieves excellent muon identification performance while maintaining exceptionally low pion misidentification rates CEPCStudyGroup:2018ghi . When combined with the signal selection strategy detailed earlier, this methodology guarantees a background-free analysis Zhang:2021orr . The signal efficiency $\epsilon_{S}$ is then evaluated from the above event selection criteria. The expected signal yield is given by $N_{S}=\mathcal{L}\times\sigma(e^{+}e^{-}\to q_{d}\bar{q}_{d})\times\epsilon_{S}$. Under the background-free hypothesis, the $2\sigma$ exclusion limit is set at $N_{S}=3$, from which we derive the exclusion contour in the $(m_{X},\kappa)$ plane, as shown in Fig. 9.
@@ VI Summary and discussion
We explored novel signatures of a leptophilic QCD-like composite asymmetric dark sector at $e^{+}e^{-}$ colliders by simulating dark parton showering, hadronization, and detector responses to map the experimental reach. The study reveals a sharp transition in the dark pion proper decay length at the muon threshold (Fig. 2). Recast constraints from monophoton searches and from slepton/stau reinterpretations limit the low-mass bi-charged scalar mediator region (Figs. 3 and 4), while combined projections for BaBar, Belle II, L-GAZELLE, and CEPC produce the characteristic band-shaped sensitivity in the $(m_{X},\kappa)$ plane as summarized in Fig. 11. After applying cosmological constraints on the $\pi_{d}$ lifetime, nearly all parameter space for $10~\text{MeV}\leq m_{\pi_{d}}<2m_{\mu}$ and $m_{X}\lesssim 10~\text{TeV}$ is excluded. Additionally, for $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, a significant fraction of parameter space with $\kappa\gtrsim 1$ and $m_{X}\lesssim 10~\text{TeV}$ is also excluded.
@@
Figure 11: Combined exclusion limits for different dark pion mass ranges. The first panel shows the exclusion region for $10~\text{MeV}<m_{\pi_{d}}<2m_{\mu}$, while the remaining three panels show benchmark points $m_{\pi_{d}}=0.32$, $0.7$ and $1~\text{GeV}$ from LHC, monophoton searches at LEP and CEPC, displaced vertex searches (BaBar, Belle II, L-GAZELLE) and DMJ search at CEPC. All four panels include the parameter space where the proper lifetime of $\pi_{d}$ exceeds one second, conservatively evaluated at $m_{\pi_{d}}=0.2~\mathrm{GeV}$ for the light mass range.
@@
The primary experimental targets we identify are (i) sparse displaced dimuon vertices at B factories (kinematics in Figs. 57, and 8, while trigger and event selection conditions are summarized in Tables 2 and 3), which Belle II can probe with realistic triggers and event selections, (ii) far detectors (L-GAZELLE) that extend sensitivity to substantially longer lifetimes, and (iii) DMJ searches at high-energy $e^{+}e^{-}$ colliders (main kinematics shown in Fig. 10 and detector assumptions in Table 6). Although our projections adopt simplifying assumptions, LFU of the mediator coupling, the benchmark relation $f_{\pi_{d}}=m_{\pi_{d}}$, and optimistic/background-free hypotheses in some regions, they nevertheless provide concrete, experimentally accessible targets and strongly motivate a coordinated program across B factories, far detectors, and Higgs factories.
@@
Finally, we outline promising directions for future work. Advanced data-scouting or trigger-level analyses for low-$p_{T}$ displaced dimuon pairs at CMS and LHCb could significantly strengthen LHC constraints LHCb:2020ysn ; CMS:2023hwl ; Born:2023vll . High-energy muon colliders could directly produce the bi-charged scalar mediator via Drell–Yan processes and thereby provide an independent probe of the parameter space. Furthermore, when the dark pion is heavier than $2m_{\tau}$, dark jets become $\tau$-rich. This regime will require dedicated strategies (e.g., displaced and boosted $\tau$-tagging, and optimized reconstruction of hadronic $\tau$ decays) to retain sensitivity.
Appendix A FCC-ee and CEPC Exclusion Region Comparison
This appendix compares the FCC-ee and CEPC experiments in terms of their exclusion regions from monophoton and DMJ searches. We adopt the same parameter settings and analysis strategies for the monophoton and DMJ signatures as in the CEPC study. The observed differences are primarily attributed to the distinct detector geometries and the modest variation in integrated luminosity between FCC-ee and CEPC experiments. The geometry and relevant transverse spatial resolutions of the FCC-ee detector are summarized in Table 7, following Refs. IDEAStudyGroup:2025gbt ; Braibant:2021wts . The transverse spatial resolutions are taken with reference to CEPC. Here we take $m_{\pi_{d}}=1.0$ GeV as an example to demonstrate this similarity-other benchmark points exhibit nearly identical behavior.
As shown in Fig. 12, we compare the parameter space exclusion capabilities between FCC-ee and CEPC for $m_{\pi_{d}}=1.0$ GeV. Considering the overall coverage of the four detector subsystems, the effective detector length of FCC-ee is slightly larger than that of CEPC, therefore FCC-ee demonstrates a mild advantage. However, this improvement is very small, and the exclusion capabilities of the two experiments are quite comparable. Therefore, for high-energy $e^{+}e^{-}$ colliders, we take CEPC as a representative example and perform detailed studies on the three BPs in the main text.
Detector Component $R_{\text{in}}$ $R_{\text{out}}$ $\sigma_{xy}$
[…]
Hadron Calorimeter $2.80\text{\,}\mathrm{m}$ $4.60\text{\,}\mathrm{m}$ $30\text{\,}\mathrm{m}\mathrm{m}$
Muon System $4.60\text{\,}\mathrm{m}$ $5.24\text{\,}\mathrm{m}$ $20\text{\,}\mathrm{m}\mathrm{m}$
Table 7: The FCC-ee detector geometry and its relevant transverse spatial resolution.
Figure 12: The comparison of exclusion limits from monophoton (left) and DMJ (right) searches between the CEPC and FCC-$e$ experiments, for a dark pion mass of $m_{\pi_{d}}=1~\text{GeV}$.